Find the absolute minimum of the function f(x, y) = 3 +xy-x-2yon the closed triangular region with vertices (1,0), (5,0), and (1,4).
1
Expert's answer
2020-10-19T15:49:21-0400

"\displaystyle\textsf{Since}\, f \, \textsf{is a polynomial, it is continuous on the closed,}\\
\textsf{bounded set}\, T,\, \textsf{where}\, T\, \textsf{is the region in the triangle.} \\

\textsf{The first order partial derivatives are} 
\\ f_x(x, y) = y - 1\, \textsf{and}\, f_y(x, y) = x - 2 \\

\textsf{Thus, the only critical points in the interior of} \,T\,\textsf{is} \,(2, 1) \\

\textsf{The boundary of}\,T \, \textsf{consists of three line segments,}\\
 L_1, L_2 \, \textsf{and} L_3. \\


\textsf{On}\, L_1,\, \textsf{we have}\, x = 1, \\

f(1, y) = 3 + y - 1 - 2y = 2 - y \\

\textsf{The function}\, f \, \textsf{is a decreasing function of}\,y \\
\textsf{along}\, L_1 \, \textsf{from}\, (1, 0) \, \textsf{to}\, (1, 4)\\

\textsf{Its maximum value is}\, f(1, 0)\, \textsf{and its minimum value is}\,f(1, 4) \\

\textsf{On}\, L_2, \, \textsf{we have} \,y = 0, \\

f(1, y) = 3 - x\\

\textsf{The function}\, f\, \textsf{is a decreasing function of}\, x \textsf{along}\\
L_2\, \textsf{from} (1, 0)\, \textsf{to}\,(5, 0) \\

\textsf{Its maximum value is}\, f(1, 0)\, \textsf{and its minimum value is}\, f(5, 0) \\


\textsf{Deriving an equation for line}\, L_3\\

\textsf{The equation can be derived}\\\textsf{from the formula for the equation}\\ \textsf{of a straight line passing through}\\
\textsf{one point with a gradient which is given by} \\

\frac{y - y_1}{x - x_1} = m(x - x_1) \\
 
\textsf{since the line is passing through}\, (1, 4) \, \textsf{and}\, (5, 0) \\

m = \frac{0 - 4}{5 - 1} = \frac{-4}{4} = -1 \\

\frac{y - 0}{x - 5} = -1 \\

\therefore y = 5 - x \\

f(x, 5 - x) = 3 + x(5 - x - 1) - 2(5 - x) \\

f(x, 5 - x) = 3 + 4x - x^2 - 10 + 2x = 6x - x^2 - 7 \\

f_x(x, 5 - x) = 6 - 2x = 0, \implies x = 3\, \textsf{and so},\, (3, 2)\, \textsf{is a critical point.} \\

\textsf{The values of the function at the}\\\textsf{critical points is given by;}\\

\begin{aligned}
f(3, 2) &= 3 + 3(2 - 1) - 2(3) = 0 \\
f(1, 0) &= 3 - 1 = 2\\
 f(5, 0) &= 3 - 5 = -2\\
f(1, 4) &= 3 + 1(4 - 1) - 2(4) = -2\\
f(2, 1) &= 3 + 2(1 - 1) - 2(1) = 1
\end{aligned} \\

\textsf{We can hence conclude that}\, -2\, \textsf{is the absolute }\\
\textsf{minimum of the function}\, f(x, y) = 3 + x(y - 1) - 2y\, \textsf{on the closed}\\
\textsf{triangular region with vertices}\, (1,0), (5, 0) \,\textsf{and}\, (1, 4)."